There are 1.5 Kgm of a gas where K = 1.3 and R = 0.38 KJ/Kgm - °R that undergo an isochoric process from p1 = 0.552 MPa, t1 = 58.5°C top2 = 1.66 MPa. During the process, there added 100 KJ of heat. Compute the heat transferred, change of internal energy and the change of entropy.


1
Expert's answer
2021-02-26T05:41:22-0500

mass of gas=1.5 Kg,K=1.3, and R=0.38 KJ/Kgm - °R and

P1 = 0.552 MPa, T1 = 58.5°C ,P2 = 1.66 MPa., Heat added to system = 100 kJ

As the process is isochoric , we know that for isochoric process work done is zero.

and heat transfer is given as , "\Delta Q= 100 kJ"

As per first law of thermodynamics

"\Delta Q= \Delta U+ \Delta W"

100 = "\Delta U +0"

So, internal energy = 100 kJ

"\Delta S= \frac{ \Delta Q}{T}= \frac{ 100 \times 1000}{ 331.5 }"

"\Delta S= 301.659 J/K" is the entropy of system